A particle executes simple harmonic motion with amplitude of 3.00cm. At what position does its speed equal half of its maximum speed?
1
Expert's answer
2021-02-11T17:11:15-0500

"x=Acos\omega t,"

"v=-A\omega sin\omega t,"

"v_{max}=A\omega,"

"v=-v_{max}sin \omega t,"

"\frac{v}{v_{max}}=-sin \omega t=\frac 12,"

"sin\omega t=-\frac 12,~~" "\Rightarrow~~" "cos\omega t=\frac{\sqrt 3}{2},"

"x_1=\frac{A\sqrt3}{2}\approx2.60~\text{cm}."